# FRIREN: Beyond Trajectories — A Spectral Lens on Time

### v1

@@ Abstract
Long-Term Time-Series Forecasting (LTSF) models often position themselves as general-purpose solutions ready to be deployed in different fields, implicitly assuming any data is pointwise predictable. Using chaotic systems like Lorenz-63 as a touchstone, we argue that geometry—not pointwise predictions—is the appropriate abstraction level for a truly dynamic-agnostic foundational model. Minimizing Wasserstein-2 distance (W2), a metric that measures geometric changes, and exposing a clear spectral view of dynamics, are key for long horizon forecasting. Our model FRIREN (Flow-inspired Representations via Interpretable Eigen-networks), implements an Augmented Normalizing-Flow block that embeds data into a normally distributed latent representation. It then produces a W2-efficient optimal path, decomposable as "rotate, stretch, rotate back, and translate". This design yields (i) locally generated, geometric structure preserving predictions that bypass underlying dynamics, and (ii) a global spectral representation that doubles as a finite Koopman operator when slightly modified, letting practitioners inspect which modes grow, damp, or oscillate across at both local and system levels. FRIREN achieves an MSE of 11.4, MAE of 1.6, and SWD of 0.96 on Lorenz-63 in a 336-in-336-out, dt=0.01 setting, decisively surpassing TimeMixer (MSE 27.3, MAE 2.8, SWD 2.1). Our model maintains effective prediction for 274 out of 336 steps, nearly 2.5 Lyapunov times. On Rossler (96-in, 336-out), FRIREN achieves an MSE 0.0349, MAE 0.0953, and SWD 0.0170, outperforming TimeMixer’s MSE 4.3988, MAE 0.886, and SWD 3.2065. It remains competitive on standard LTSF datasets such as ETT and Weather. FRIREN bridges the gap between modern generative flows and classical spectral analysis, making long-term forecasting both accurate and interpretable—setting a new standard for LTSF model design.
@@ 1 Introduction
Deep learning models, including transformers, linear methods, and LLMs (Wu et al., 2022; Zhou et al., 2021; Zeng et al., 2023; Nie et al., 2022; Wang et al., 2024; Das et al., 2024), are widely applied to Long-Term Time Series Forecasting (LTSF). Here, past values in a time-delayed embeddings $x\_1,\dots,x\_n$ map into future values $y\_1,\dots,y\_k$ (with $x_{n+1}\coloneqq y_{1}$), typically over horizons of 96, 196, 336, or 720. Despite rapid advances, the field lacks a clear, systematic understanding of predictability itself. All data-driven methods share a core obstacle: we only have purely observational sequences—no interventions and an unknown measurement error—so attaining genuine causal insight is elusive. One approach, championed by Brunton et al. (2016b) and Conti et al. (2024), seeks to approximate governing equations by selecting a sparse set of functions from a predefined dictionary. This approach is great for deterministic systems where a latent differential equation exists. The interpretability gives us true insights into the causal mechanisms. However, for real-world predictions where random nonstationarity and stochastic factors exist, the difficulty grows tremendously.
The LTSF community, by contrast, adopts a purely data-driven stance, positioning models as dynamics agnostic, ready to handle deterministic, stochastic dynamics, as well as additional nonstationary shocks. Modules designed for natural language semantic alignment, periodic patterns, smoothed trends, and noise damping handle the data. Implicitly, it is assumed that data is simply predictable, although in a black-box and data-driven way. This unwarranted optimism received sharp critique in Bergmeir (2023, 2024), where simple last-value predictors outperform state-of-the-art models on nearly random-walk data. This indicates predictability analysis is essential. Zeng et al. (2023) also famously illustrated the problem of blind faith in transformer modules. For a long time, the extraordinary long-range alignment ability of the transformer model (Vaswani et al., 2017) on other domains is implicitly treated as a source of predictability. At the current stage, despite tremendous progress in recent years, a sharp yet fair assessment expressed by many in the community is that patching (as in PatchTST (Nie et al., 2022)) and smoothing (as in DLinear (Zeng et al., 2023) and TimeMixer (Wang et al., 2024)) are what “really works”.
Predicting Chaos: Can You Predict the Butterfly?
We want to reexamine the problem of predictability for a purely data-driven, dynamic-agnostic model that is designed to be foundational and ready to be deployed in different fields. Here is a simple challenge: can you predict the Butterfly? This refers to the 1963 Lorenz system (Lorenz, 1963) (see Figure 3 and for visualization and Section A.2.1 for governing equations), the famous “butterfly effect” where metaphorically a flap of a butterfly’s wings in Brazil can set off a tornado in Texas (Lorenz, 1972). Three simple equations produce extraordinarily complex behaviour. Like weather forecasting where it originated, the Lorenz system shows patterns that are predictable in the short-term but complex and unpredictable long-term. Its butterfly-shaped attractor, i.e. a set of states that a system evolves towards, illustrates the mesmerizing dance of particles between “wings”: they swirl within one lobe before irregularly switching to the other. Our challenge is specific: can LTSF models predict the Lorenz system for the next 96, 196, 336, and 720 steps ahead, where each step advances the system by 0.01 model time units (MTU)? Or for a much longer time horizon?
Long-Term Forecasting and the Challenge of Chaos
The Lorenz system is a canonical example of chaotic system where a purely deterministic process exhibits random-like behaviour. Over sufficiently long horizons, the system becomes unpredictable, due to the defining property of the chaotic system, the sensitive dependence on initial conditions: “any two points that lie arbitrarily close to each other will move apart so dramatically that it is impossible to verify, after a reasonably short period, whether they were ever close to each other” (Osinga, 2018). Remarkably, even with knowledge of the exact equations and initial conditions, tiny measurement errors in them grow exponentially, making long-term prediction impossible even in such ideal cases.
Within a reasonably short horizon such as what we proposed, the Lorenz system remains especially challenging. For autoregressive models that attempt to infer trajectories step by step, while there are recent breakthroughs with mixture model such as Wang and Li (2024) (blending data-driven prediction and data-driven equation discovery), pure data-driven approach such as LSTM (Hochreiter and Schmidhuber, 1997) still struggles. Details can be found in (Wang and Li, 2024). The reason is that in the Lorenz system, trajectories will never exactly repeat themselves or self-cross while covering a thin shell-like attractor surface. Such fundamental irregularity makes stepwise patterns misleading. Compounding prediction error quickly leads to inevitable divergence from the true trajectory.
Geometry over Trajectory
In the Lorenz system, chaos’s hidden order is its attractor. A remarkable mathematical discovery, the Takens’ Embedding Theorem (Takens, 1981) (see Theorem A.1 for details), reveals that under appropriate conditions, a sufficiently long time-delay embedding of a single variable contains all the necessary information to reconstruct the system’s attractor. This reconstruction, while not identical in appearance to the butterfly-shaped form, is related to it through a diffeomorphism—a smooth, invertible transformation that preserves the essential topological structure and dynamics of the original system. Intuitively, it means the butterfly-shaped attractor is bent or stretched, but not glued and torn.
This attractor represents the collective distribution of paths that a particle can take as it repeatedly returns to a neighborhood, capturing the full range of possible trajectories. It is a clear, stable, and predictable geometric object. While exact trajectories remain unpredictable, once a prediction is placed along this deformed butterfly-shaped attractor, the range of error is well controlled. The answer to the question of Can you Predict the Butterfly? is “No, but if we can reconstruct the theoretically guaranteed attractor and place our prediction accordingly, we cannot be too wrong within a reasonable window”.
Rethinking Prediction: At What Level of Abstraction
Chaotic behaviour isn’t niche; it’s ubiquitous across domains where accurate forecasting is most crucial: climate modelling, power grids, neuronal activity, financial markets, and epidemiology. Most revealingly, even standard LTSF benchmarks such as the ETT dataset (Zhou et al. (2021)) have been proven to exhibit chaotic characteristics ((Hu et al., 2024), Section E.2). Therefore, the Lorenz system is therefore not merely an academic curiosity—it represents a stress test on a fundamental class of dynamics pervading real-world systems. A truly dynamic-agnostic LTSF model—one that claims general applicability—must necessarily demonstrate competence on such systems. LABEL:app:tab:lorenz63 illustrates this point dramatically: DLinear (Zeng et al., 2023), a simple and effective model that performs competitively against specialized model such as TimeMixer (Wang et al., 2024) and PatchTST (Nie et al., 2022) on standard benchmark such as ETTm2 (see LABEL:app:data:ettm2 for comparisons), shows 50$\times$ higher error than our model FRIREN, and 20$\times$ higher error than the other two, on a 96-in-96-out short-term Lorenz prediction task. Notably, this performance gap narrows sharply when the prediction horizon is expanded to 720, highlighting the fundamental challenges of long-term prediction in chaotic systems.
Prediction is hard. On one extreme, a purely random walk is fundamentally unpredictable; similarly (but to a lesser degree), a stable system subjected to persistent, unknown shocks becomes difficult to predict. On the opposite extreme, a purely chaotic system is inherently unpredictable over the long term; similarly (but less severely), a periodic system that occasionally shifts into chaotic regimes also poses significant prediction challenges. In real-world forecasting, difficult tasks often lie between these extremes. A purely data-driven foundational LTSF model intended for real-world deployment should be ready to handle both kinds of dynamics by identifying what remains stable across these varied conditions. For this purpose, we argue that the most robust abstraction level is pattern matching via geometry. If a pattern is conserved, there must be some mechanism that conserves the order amid chaos. Stable geometry is a credible manifestation of true mechanisms. An ideal model simultaneously recognizes genuine geometry changes and resists perturbation, a designing principle we termed geometric conservatism. This geometric approach can serve as a robust, first-line diagnostic tool when the underlying dynamics remain unknown; the more ambitious goal of governing equation inferring (difficult under nonstationary shocks) or exact trajectory prediction (difficult for chaotic dynamics with mere data-driven methods) can be achieved with specialized approaches once we have significant prior knowledge about the system.
@@ 2 Method: Geometric Prediction via Location-Scale Optimal Transport
Mean squared error (MSE), defined as $\mathbb{E}_{x}[\mathbb{E}_{y\sim P(y|x)}[\|y_{pred}(x)-y_{real}\|^{2}]]$, measures only the average pointwise accuracy, ignoring the superior distributional fidelity of the first oracle’s prediction. Similarly, it fails to capture the intuitive similarity in the second example. The key element lost in MSE is invariance to phase shifts.
@@ 3 Experiments
In terms of scalability, FRIREN is efficient. It avoids expensive components, and the Householder rotations are lightweight. The primary computational cost arises from spectral normalization, which can slow down training but is essential for stability on some datasets. Finally, Our complete codebase is available at https://anonymous.4open.science/r/LTSF_model-C6B8/ for reproducibility.

### v3

@@ Abstract
Long-Term Time-Series Forecasting (LTSF) models often position themselves as general-purpose solutions ready to be deployed in different fields, implicitly assuming any data is pointwise predictable. Using chaotic systems like Lorenz-63 as a touchstone, we argue that geometry—not pointwise predictions—is the appropriate abstraction level for a truly dynamic-agnostic foundational model. Minimizing Wasserstein-2 distance (W2), a metric that measures geometric changes, and exposing a clear spectral view of dynamics, are key for long horizon forecasting. Our model FRIREN (Flow-inspired Representations via Interpretable Eigen-networks), implements an Augmented Normalizing-Flow block that embeds data into a normally distributed latent representation. It then produces a W2-efficient optimal path, decomposable as "rotate, stretch, rotate back, and translate". This design yields (i) locally generated, geometric structure preserving predictions that bypass underlying dynamics, and (ii) a global spectral representation that doubles as a finite Koopman operator when slightly modified, letting practitioners inspect which modes grow, damp, or oscillate across at both local and system levels. FRIREN achieves an MSE of 11.4, MAE of 1.6, and SWD of 0.96 on Lorenz-63 in a 336-in-336-out, dt=0.01 setting, decisively surpassing TimeMixer (MSE 27.3, MAE 2.8, SWD 2.1). Our model maintains effective prediction for 274 out of 336 steps, nearly 2.5 Lyapunov times. On Rossler (96-in, 336-out), FRIREN achieves an MSE 0.0349, MAE 0.0953, and SWD 0.0170, outperforming TimeMixer‘s MSE 4.3988, MAE 0.886, and SWD 3.2065. It remains competitive on standard LTSF datasets such as ETT and Weather. FRIREN bridges the gap between modern generative flows and classical spectral analysis, making long-term forecasting both accurate and interpretable—setting a new standard for LTSF model design.
@@ 1 Introduction
Deep learning models, including transformers, linear methods, and LLMs (Wu et al., 2022; Zhou et al., 2021; Zeng et al., 2023; Nie et al., 2022; Wang et al., 2024; Das et al., 2024), are now standard for Long-Term Time Series Forecasting (LTSF). The community largely treats them as dynamics-agnostic tools, implicitly assuming that complex patterns are simply pointwise predictable in a data-driven way—even under a simultaneous mix of stochastic, chaotic and non-stationary influences. This demand for robustness is justified, as mixtures of these dynamics are ubiquitous in real-world systems, from climate and power grids to finance. Even standard LTSF benchmarks like the ETT dataset (Zhou et al., 2021) exhibits known chaotic characteristics ((Hu et al., 2024), Section E.2). However, a strong solution in a very general setting cannot just be assumed to exist thanks to the successful transformer architecture Vaswani et al. (2017) without careful analysis of the predictability of datasets. In sharp critiques, Bergmeir (2023) questions benchmarking on nearly random-walk dataset (Exchange Rate), and Zeng et al. (2023) demonstrated transformer modules alone, without the widely acknowledged technique such as patching and smoothing, offer limited accuracy improvement against a simple linear model (DLinear) where one regression predicts the smoothed trend and another predicts the residual.
This paper challenges the dynamics-agnostic assumption with a canonical problem: Can you predict the butterfly? This refers to the 1963 Lorenz system (Lorenz, 1963) (Figure 3 for visualization and Section A.2.1 for governing equations), whose butterfly-shaped attractor (a set of states that a system evolves towards) is the emblem of the "butterfly effect". As a purely deterministic process, its defining property is a sensitive dependence on initial conditions: “any two points that lie arbitrarily close to each other will move apart so dramatically that it is impossible to verify, after a reasonably short period, whether they were ever close to each other” (Osinga, 2018). Even if the governing equations were known and initial conditions measured near perfectly (both are very hard to reverse engineer), any infinitesimally small errors in measurement still grow exponentially, rendering long-term trajectory forecasting impossible, even in such ideal situations. For purely data-driven autoregressive models, the challenge is even greater: the system’s non-repeating, non-crossing trajectories make learned stepwise patterns misleading, causing prediction errors to compound catastrophically.
Amid this chaos, the hidden order is the system’s geometry. Takens’ Embedding Theorem (Takens, 1981) provides a remarkable guarantee: a time-delay embedding of a single variable can reconstruct a diffeomorphism of the system’s full attractor, preserving its topological structure (see Theorem A.1). A diffeomorphic transformation is a smooth and invertible transformation; intuitively, it means the butterfly-shaped attractor is bent or stretched, but not glued and torn. This stable, predictable geometric object captures the complete distribution of possible system states. Thus, the answer to our question is: “No, you cannot predict the exact trajectory. However, if you can reconstruct the attractor, you can place your prediction on it, and the range of error will be well-controlled." Thus, our core claim: predict geometry, not trajectory.
In the Lorenz system, the difficulty arises from the mesmerizing dance of particles between ‘wings’: they swirl within one lobe before irregularly switching to the other. Therefore, asking LTSF models if they can predict the next 96, 196, 336, and 720 steps ahead, where each step advances the system by 0.01 model time units (MTU), is a crucial stress test. Our results in LABEL:app:tab:lorenz63 are stark: on a short-term (96-step) forecast, DLinear—a strong benchmark performer on ETT—has 50$\times$ higher error than our model, FRIREN, and 20$\times$ higher than PatchTST and TimeMixer. This demonstrates that models optimized for stochastic trends can fail dramatically on deterministic chaos.
Forecasting is a trade-off between task strength and structural generality. Enhanced with governing equation discovery techniques (e.g. Brunton et al. (2016b) and Conti et al. (2024)), even autoregressive model exclusively designed for chaotic system can achieve a remarkable stability (Wang and Li, 2024). A truly general model, however, must be robust to different sources of unpredictability: (1) stochasticity, (2) deterministic chaos, and (3) external non-stationary shocks. We propose that the most robust abstraction level for such foundational model is pattern matching via geometry. Our model, FRIREN, embodies the principle of geometric conservatism: it aims to identify and preserve the stable geometric structures that represent conserved mechanisms amidst noise and chaos. By focusing on learning geometry-preserving maps, it excels where trajectory-based models fail, offering a new path toward building more robust and interpretable forecasting systems.
@@ 2 Method: Geometric Prediction via Location-Scale Optimal Transport
Recall that in LTSF, the task is a direct multi-step prediction where past values in a time-delayed embeddings $x_{1},\dots,x_{n}$ map to future values $y\_1,\dots,y\_k$ (with $x_{n+1}\coloneqq y_{1}$), typically over horizons of 96, 196, 336, or 720. Mean squared error (MSE), defined as $\mathbb{E}_{x}[\mathbb{E}_{y\sim P(y|x)}[\|y_{pred}(x)-y_{real}\|^{2}]]$, thus measures only the average performance, ignoring the superior distributional fidelity of the first oracle’s prediction and failing to capture the intuitive similarity in the second example. The key element lost in MSE is invariance to phase shifts.
@@ 3 Experiments
In terms of scalability, FRIREN is efficient. It avoids expensive components, and the Householder rotations are lightweight. The primary computational cost arises from spectral normalization, which can slow down training but is essential for stability on some datasets. Finally, Our complete codebase is available at https://anonymous.4open.science/r/LTSF_model-03BB/ for reproducibility.